A circular loop of 80 turns with radius 15.0 cm is perpendicular to a magnetic field of 0.52 T. How much induced voltage is created if it is rotated in 0.20 s so that its face is parallel to the field?





1
Expert's answer
2021-12-08T10:30:43-0500

Induced"EMF = \frac{Bπr^2n}{t}"

Where B=magnetic field=0.52T

r=radius=15cm=0.15m

n=no. Turns = 80

t= time=20sec

"EMF = \frac{0.52×π×0.15^2×80}{0.2}=14.7v"


EMF=14.7V